OpenStudy (anonymous):

convert 11010 to base ten

OpenStudy (anonymous):

= 2^4 + 2^3 + 0*2^2 + 2^1 + 0*2^0 = 2^4 + 2^3 + 2 = 16 + 8 + 2 = 26